OpenStudy (anonymous):

Solve 3x^2=-30x+24? I know that I need to use the quadratic formula, but I am really lost... :/

jimthompson5910 (jim_thompson5910):

3x^2=-30x+24 3x^2+30x-24 = 0 3(x^2 + 10x - 8) = 0 x^2 + 10x - 8 = 0/3 x^2 + 10x - 8 = 0 a = 1, b = 10, c = -8 \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(10)\pm\sqrt{(10)^2-4(1)(-8)}}{2(1)}\] \[\Large x = \frac{-10\pm\sqrt{100-(-32)}}{2}\] \[\Large x = \frac{-10\pm\sqrt{100+32}}{2}\] I'll let you finish. Remember to simplify as much as you can.

OpenStudy (anonymous):

When dividing, I divide both -10 and the sqrt of 132 by 2, correct?

jimthompson5910 (jim_thompson5910):

well you first simplify sqrt(132) to get sqrt(132) sqrt(4*33) sqrt(4)*sqrt(33) 2*sqrt(33) So sqrt(132) simplifies to 2*sqrt(33)

jimthompson5910 (jim_thompson5910):

then you would divide both -10 and 2*sqrt(33) by 2

jimthompson5910 (jim_thompson5910):

So your answers are \[\Large x = -5+\sqrt{33} \ \text{or} \ x = -5-\sqrt{33}\] which you can write them as \[\Large x = -5 \pm \sqrt{33}\]

OpenStudy (anonymous):

Thanks, I wasn't sure how to write it out. When I did it on my own I think I was missing steps and not completing the whole process :)

jimthompson5910 (jim_thompson5910):

you're welcome